OpenStudy (anonymous):

how to solve these: 1. 2x+y-z=2 x+2y+4z=1 5x+y-7z=4 2. 3/x + 4/y + 6/z = 1 9/x + 8/y - 12/z = 3 9/x - 4/y + 12/z = 4 3. x+y+z=0 y+z-2w=-5 2x-y+3w=9 x+3z-w=-1 4. 2x-y+z-w =5 2y-z+2w=-4 x+y-3w=0 3y+2z-2w=1 by elimination and substitution.. help me please ..

terenzreignz (terenzreignz):

This is a nightmare :3

terenzreignz (terenzreignz):

I reckon the first is the easiest, so let's get down to business...

OpenStudy (austinl):

Oh my.

OpenStudy (austinl):

FUN!!!!1!!!!

terenzreignz (terenzreignz):

I shudder at what you find fun. Now, as for the first set of equations, the most straightforward way to go about this is substitution...

OpenStudy (anonymous):

huhu T_T .. hate it ..

terenzreignz (terenzreignz):

Let's look at the first equation of the first system... 2x + y - z = 2

terenzreignz (terenzreignz):

From here, it follows that 2x + y - 2 = z correct?

OpenStudy (anonymous):

yes? haha .. not sure. hm, ah! yeah.. correct!

terenzreignz (terenzreignz):

I don't like 'not sure'. Be sure, or clarify with me.

terenzreignz (terenzreignz):

What didn't you get?

OpenStudy (primeralph):

Someday, but not today.

OpenStudy (anonymous):

ahh.. i know na . :) next?

terenzreignz (terenzreignz):

z = 2x + y - 2 Now, in the other two equations x + 2y + 4z = 1 5x + y - 7z = 4

terenzreignz (terenzreignz):

I want you to replace z with that new expression. Tell me what you get.

OpenStudy (anonymous):

x + 2y + 4 (2x+y-z) = 1 5x + y - 7 (2x+y-z) =4 right?

terenzreignz (terenzreignz):

No.

OpenStudy (anonymous):

ee how?

terenzreignz (terenzreignz):

\[\huge \color{red}{z= 2x + y - \boxed2}\]

OpenStudy (anonymous):

ahh.. i know na .. x + 2y + 4 (2x+y-2) = 1 5x + y - 7 (2x+y-2) =4 right?

terenzreignz (terenzreignz):

Now, simplify.

OpenStudy (anonymous):

x + 2y + 8x + 4y - 8 = 1 5x + y - 14x - 7y + 14 = 4

terenzreignz (terenzreignz):

...now combine like terms

OpenStudy (texaschic101):

wouldn't elimination be easier on the 1st one

OpenStudy (primeralph):

And out of nowhere,a pretty @texaschic101 appears.

terenzreignz (terenzreignz):

I don't know which is easier.

OpenStudy (anonymous):

9x + 6y = 9 -9x - 6y = -10

OpenStudy (texaschic101):

lol...hello primeralph

OpenStudy (primeralph):